OpenStudy (anonymous):

The base of a ten-foot ladder is being pushed away from the wall it is leaning against . If the base is moving at a rate of 2 feet per minute find the rate at which the top of the ladder is moving when the lower end of the ladder is 3 feet from the wall describe the significance of the sign of your results . discs s whether or not this sign depends on which side of the wall the ladder is leaning .

OpenStudy (anonymous):

did you draw a picture?

OpenStudy (anonymous):



OpenStudy (anonymous):

okay let me work it out

OpenStudy (anonymous):



OpenStudy (anonymous):

no, h is not changing with time. (h is fixed at 10 ft., the length of the ladder).

OpenStudy (anonymous):

2x*dx/dt=100

OpenStudy (anonymous):

dh/dt = 0

OpenStudy (anonymous):

\[x^2+y^2=h^2 \Rightarrow 2x \frac{ dx }{ dt }+2y \frac{ dy }{ dt }=0\]

OpenStudy (anonymous):

they're asking for dy/dt

OpenStudy (anonymous):

oh, so constant is 10 and the derivative of it is 0

OpenStudy (anonymous):

how come the y isn't a constant?

OpenStudy (anonymous):

look at the triangle... as x increases, y has to decrease because h stays constant.

OpenStudy (anonymous):

oh... will y be sqrt 91 then?

OpenStudy (anonymous):

yes... y = sqrt(100-9) = sqrt(91)

OpenStudy (anonymous):

its a negative sign because the y is shrinking right?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

if the base is moving away from the wall then x is always increasing and subsequently, y will always be decreasing.

OpenStudy (anonymous):

if the base moves towards the wall then x will be decreasing and in that case, y would be increasing.

OpenStudy (anonymous):

oh~~~ I get it now xD

OpenStudy (anonymous):

just get a 10' ladder and try the experiment... maybe use a strobe or slow motion video.

OpenStudy (anonymous):

Alright! Thanks a lot!

OpenStudy (anonymous):

you're welcome!